OpenStudy (anonymous):

The formula a=46c gives the floor area (a) in square meters that can be wired using (c) circuits a. Solve a=46c for c b. If a room is 322 square meters, how many circuits are required to wire this room?

OpenStudy (anonymous):

a) c = a/46

OpenStudy (anonymous):

From c = a/46 we calculate c Value given a = 322 square meters. c = 322/6 = 5.3 = 5 (Approx)